# GmNet: Revisiting Gating Mechanisms From A Frequency View

### v1

@@ Abstract
Gating mechanisms have emerged as an effective strategy integrated into model designs beyond recurrent neural networks for addressing long-range dependency problems. In a broad understanding, it provides adaptive control over the information flow while maintaining computational efficiency. However, there is a lack of theoretical analysis on how the gating mechanism works in neural networks. In this paper, inspired by the convolution theorem, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We investigate the interact between the element-wise product and activation functions in managing the responses to different frequency components. Leveraging these insights, we propose a Gating Mechanism Network (GmNet), a lightweight model designed to efficiently utilize the information of various frequency components. It minimizes the low-frequency bias present in existing lightweight models. GmNet achieves impressive performance in terms of both effectiveness and efficiency in the image classification task. Our code is public at https://github.com/YFWang1999/GmNet
@@ 1 Introduction
The gating mechanism has been widely used in recurrent neural networks (RNNs) [13, 3] to address long-term dependency issues. With the flourish of Transformers [29], self-attention-based architectures have largely replaced RNNs as the core model in natural language processing (NLP). Despite their success, scaling Transformers to long sequences remains challenging due to the quadratic complexity of global attention. Linear gating mechanisms (i.e. Gated Linear Units [4]) have been rediscovered as an effective approach to address this problem.
Gated Linear Units (GLUs) introduce a gating mechanism that adaptively controls information flow, allowing the model to focus on relevant features and ignore less important ones. These architectures are simple to implement, and have no apparent computational drawbacks. Therefore, models built with GLUs have been developed in various tasks across NLP [5, 6, 10] and Computer Vision [16, 18, 33].
@@
While architectures based on linear gating mechanisms have demonstrated the effectiveness, their theoretical analysis have not received enough attention. Except intuitive explanations, a deeper examination of their mathematical properties could provide valuable insights for the development of efficient gating mechanisms. Inspired by the Fourier interpretation of element-wise product, we propose to investigate training dynamics of neural networks (NNs) with gating mechanisms from a frequency perspective. This approach helps us understand how frequency components are selectively amplified or suppressed during learning. By the frequency analysis, we aim to uncover how GLUs influence the spectral properties of neural networks and affect the overall performance.
@@
In this paper, we systematically explore how gating mechanisms impact the training dynamics. We also provide a comprehensive understanding of how the element-wise product and activation functions interplay to shape the frequency response of neural networks. The first observation is inspired by the convolution theorem and convolution operation in the frequency domain. It indicates that element-wise multiplication in the spatial domain, equivalent to convolution in the frequency domain, inherently introduces new frequency components. It may enhance the learned representations for various frequency domains. However, as demonstrated in prior studies [31, 34], NNs are sensitive to high-frequency noise, which may adversely influence training stability and generalization. Therefore, it is necessary to include activation functions within the gating units to regulate the newly introduced frequency components. Activation functions can broadly be classified based on their differential properties. Smooth activation functions (GELU [12] and Swish), characterized by continuous higher-order derivatives, tend to introduce fewer high-frequency components. On the other hand, activation functions (ReLU) with discontinuous higher-order derivatives are more likely to encourage the learning of high-frequency components.
@@
Benefited by the sufficient study from the frequency view, we find that the existing lightweight methods are not aware of how to efficiently and effectively utilize the high-frequency information. In this paper, we incorporate the simplest gated linear unit into the current block design and come up with a competitive lightweight model named as Gating Mechanism Network (GmNet). Our proposed network can more effectively capture information across different frequencies, offering both theoretical and practical advantages. With a simplified architecture of the network design, GmNet is able to achieve superior results compared with existing efficient modals without training with advanced techniques. For example, compared to EfficientFormer-L1 [14], GmNet-S3 outperforms $4.0\%$ on top-1 accuracy on ImageNet-1K with 4x faster latency on an A100 GPU.
We succinctly summarize and emphasize the key contributions of this work as following points:
@@
  • We present a comprehensive study on how gated linear units influence neural network training dynamics from a frequency perspective. Through both theoretical analysis and experimental evidences, we demonstrate how the interaction between element-wise multiplication and activation functions shapes the frequency response of NNs.
  • Inspired by the study of the gating mechanism, we proposed a competitive lightweight model, named Gating Mechanism Network (GmNet), which achieves promising performance without the need for strong training strategy, surpassing numerous efficient designs.
@@ 2 Related Work
Gated Linear Units. Gating mechanisms have been extensively utilized in neural networks for sequence data processing tasks [13, 7, 9, 4]. With the rise of Transformer architectures, the Gated Linear Unit [24] has been revisited and recognized as an efficient and effective enhancement for modern deep learning models. Recent research has increasingly focused on the integration of gating mechanisms into diverse network architectures of different tasks, including gate-MLP [16], Mamba [10] and Llama3 [6]. While many existing studies provide high-level insights into the functional role of gating mechanisms, there is a notable lack of in-depth analysis of their underlying learning processes. In this work, we conduct a series of experiments combined with solid theoretical analysis to elucidate the roles of gating mechanisms in enhancing the learning of neural networks. Specifically, we investigate the types of information that gating encourages networks to capture from a frequency perspective.
@@
Frequency Learning. Understanding the learning dynamics of neural networks from a frequency perspective is an increasingly interesting topic. It has been found that neural networks respond differently to frequency information depending on the tasks. It has been investigated that neural networks tend to learn low-frequency components quickly before fitting the high-frequency ones for regression tasks [34, 25, 22]. In classification tasks, by investigating how neural networks respond to different frequency components during training, researchers found that there exists a trade-off between the accuracy and robustness [31]. NNs can improve the accuracy by learning more from high frequency components but they are also sensitive to high frequency noise which may disturb the training of NNs. In this paper, we investigate how GLUs optimize the trade-off which drive us to propose a new efficient model design.
@@
Lightweight Networks. The existing lightweight networks can be categorized into two types: (1) Pure convolution-based networks like MobileOne [28], StarNet[18] and RepVit[30]; (2) Self-attention-based architectures including EfficientMod[19] and EfficientFormerV2[14]. Although existing network designs already achieve promising performance, they still exhibit significant shortcomings from the perspective of the frequency view. Both convolution and self-attention are learning methods with a low-frequency bias [26, 1]. Furthermore, because compact networks typically focus on minimizing parameters and simplifying model structure, their capacity to capture high-frequency details may be restricted. In this paper, we propose a new competitive and efficient network that demonstrates the effectiveness of learning from higher frequency components.
Figure 2: Block design of different variants of ResNet18 where $\odot$ represents the element-wise product and $\sigma$ means the activation function. In this paper, we primarily investigate GLUs using ReLU and GELU as activation functions where corresponding variants can be named as Res18-Gate-ReLU or Res18-Gate-GELU.
@@ 3 Revisiting Gating Mechanisms from A Frequency View
To better understand how gating mechanisms work, this paper investigates the behavior of a convolutional neural network (CNN) in classification tasks from a frequency perspective. We begin by defining the components associated with different frequency bands and outlining the details of our experimental setup.
With decomposing the raw data into a series data $\mathbf{z}=\{\mathbf{z}_{i}\}_{i=0}^{n}$ , $\mathbf{z}_{i}$ represent different frequency components of an image $\mathbf{x}$. Denoting a series thresholds $R=[0,r_{1},\dots,r_{N},+\infty]$, we have the following equations:
$\mathbf{z}=\mathcal{F}(\mathbf{x}),\quad\mathbf{z}_{0},\dots,\mathbf{z}_{n}=\theta(\mathbf{z};R)$ (1)
where $\mathbf{z}_{i}=\mathcal{F}\left(\mathbf{x}_{i}\right)$ is the 2D Discrete Fourier Transform of $\mathbf{x}_{i}$. $\theta(\cdot;r)$ denotes a thresholding function that separates the different frequency components from $\mathbf{z}_{i}$ according to the radii $r_{1},\dots,r_{N}$. Formally, for an one-channel image $x\in\mathcal{N}$ where $\mathcal{N}$ possible pixel values. Then, we obtain $\mathbf{z}\in\mathbb{C}^{n\times n}$, where $\mathbb{C}$ denotes the set of complex numbers. The value of $\mathbf{z}$ at coordinates $(x,y)$ is denoted by $\mathbf{z}(x,y)$, while $(c,c^{\prime})$ represents the centroid. The formal definition of $\mathbf{z}_{i}$ is given as:
$\mathbf{z}_{i}=\{\mathbf{z}(x,y)\mid r_{i-1}\leq d\left((x,y),(c,c^{\prime})\right)\leq r_{i}\},$
we consider $d(\cdot,\cdot)$ as the Euclidean distance, which operates independently on every channel of pixels.
All experiments in this section utilize ResNet18 [11] as the backbone, implemented in PyTorch [21]. Modifications to the network blocks are depicted in Fig. 2. The network is trained on the CIFAR-10 dataset using CrossEntropyLoss, with the SGD optimizer set to a learning rate of $0.1$. We evaluate the classification performance on different frequency components of the input images at each training epoch. Changes in accuracy over time provide insights into the learning dynamics within the frequency domain [31]. To avoid the occasionality, we calculated the average over three training runs.
(a) Comparison between Res18 and Res18-Ewp.
(b) Comparison between Res18-Gate-ReLU and Res18-Ewp.
(c) Comparison between Res18-Gate-ReLU and Res18-Gate-GELU.
(d) Comparison among Res18, Res18-Ewp. Res18-Gate-ReLU and Res18-Gate-GELU.
Figure 3: Learning curves of Resnet18 and its variants for $100$ epochs, together plotted with the accuracy of different frequency components $\mathbf{z}_{i}$. We set $\{r_{1},r_{2},r_{3}\}$ to $\{6,12,18\}$. All curves of $\mathbf{z}_{i}$ are from the test set. The legends can be found in the top of Fig. 3(a). All the variants shows almost same training curves and very close testing curves of the raw data where the difference is lower than $0.03$ on average. In this case, a difference value of $0.1$ can be seen as a significant change. When accuracy falls below $15\%$, it can be considered close to the random selection, suggesting that neural networks struggle to learn from this particular frequency range. For a better view, we only put curves of two variants in a single image for figures a, b and c. Also, we plot all curves in a single plot and highlighted the most upper curves for each images in (d) to have a comprehensive understanding. More results on other settings can be found in the supplementary.
@@ 3.1 Effect of Element-wise Product
To verify this insight, we conduct experiments on the Res18 and Res18-Ewp to find the effect of element-wise product. As shown in Fig. 3(a), the accuracy curve of Res18-Ewp visibly surpass that of Res18 for decomposed parts in ranges of $(0,r_{1})$ and early epochs in ranges of $(r_{1},r_{2})$. The results demonstrate that broader spectral ranges enable the network to learn faster and better across multiple frequency components. Empirically, it is usually the lower-frequency components. However, examining the accuracy curve for components within the range $(r_{2},r_{3})$ reveals a significant drop in accuracy growth after $40$ epochs compared to the base model. It indicates that although element-wise product can encourage the model learn faster on some frequency components, newly introduced high-frequencies can become the noise to disturb the effective learning on high-frequencies. Therefore, the activation layer is necessary to regulate the newly generated frequencies, minimizing the impact of high-frequency noise.
@@ 3.2 How Activation Function Works?
In this paper, we mainly focus on two kinds of activation functions based on their properties.
  • Smooth activation: Smooth function means the differential coefficient is continues like GELU and Swish. This type of functions usually tends to produce lower frequency components.
  • Non-smooth activation: Functions like ReLU and ReLU6 have discontinuous derivatives, which can introduce more high-frequency components in gating mechanisms.
Specially, we select GELU and ReLU as representations to reveal how two kinds of activation functions control the new generated frequencies differently.
3.2.1 Theoretical Demonstrations
We firstly give a brief mathematical analysis of how the differential properties of activating functions affect the frequency extension in GLUs.
@@ 3.2.1 Theoretical Demonstrations
where $f^{(n)}(t)$ represents the $n$-th derivative of $f(t)$. The proof of Lemma 2 can be easily obtained by the mathematical induction based on Theorem 2. From Eq. 2, we can find that $|F(\omega)|\leq\frac{C}{|\omega|^{n}}$ where $C$ is a sufficiently large constant. It indicates that if higher-order derivatives exist and are continuous, the high-frequency components of the Fourier transform decay more quickly with a rate of $|\omega|^{n}$. Conversely, if higher-order derivatives do not exist or are discontinuous, the decay of the high-frequency components of the Fourier transform slows down. Notably, for the smooth activation like $\operatorname{GELU}(x)=x\cdot\frac{1}{2}[1+\operatorname{erf}(x/\sqrt{2})]$, it is infinitely differentiable across the entire real domain, with all orders of derivatives being continuous and finite. The continuity and finite values of higher-order derivatives cause the high-frequency components of the Fourier transform to decay quickly, resulting in fewer high-frequency components. On the other hand, discontinuity of higher-order derivatives can lead to generating more high-frequencies. Non-smooth activation may encourage remaining more higher frequency components in the gating mechanism.
@@
Figure 4: The illustration of GmNet architecture. GmNet adopts a traditional hybrid architecture, utilizing convolutional layers to down-sample the resolution and double the number of channels at each stage. Several blocks are repeated to extract features.
3.2.2 Experimental Analysis
As shown in Fig. 3(b), compared to the model without activation, the complete gating unit (using ReLU) learns better in higher frequency components compared to Res18-Ewp, like in $\mathbf{z}_{i}\in(r_{2},r_{3})$. It effectively avoids the disturbance of high-frequency noise. Moreover, we have conducted contrast experiments of using different kinds of activation functions. From Fig. 3(c), it is obvious that the model using non-smooth activation (ReLU) performs better in learning from the higher-frequencies like $(r_{2},r_{3})$. The superiority of the variant with GELU when $r\in(0,r_{2})$ shows the capability of the smooth activation function in learning lower-frequency components. Also, the performance of Res18-Gate-GELU on $(r_{2},r_{3})$ which is under $20\%$ also indicates the low-frequency bias for the smooth activation function. Additionally, as shown in Fig. 3(d), GELU and ReLU demonstrate their superiority among the variants in capturing low-frequency and high-frequency components, respectively. It indicates that the model with a complete gating unit can better capture the information for various ranges of frequency based on their properties.
Discussions. Different datasets and training strategies may lead to different dynamics which has been addressed in [31, 26]. For the small dataset, the convergence is faster which means NNs can easily capture some certain frequency components and then learning from other frequency components becomes very slow. Therefore, the generated high-frequency noise has a greater impact on disturbing the learning from high-frequencies. We will show more evidences and discussions of the difference between various activation functions in ablation studies on a larger dataset. Moreover, from the performance on very high-frequency components (for $\mathbf{z}_{i}\in(r_{3},+\infty)$), it is evident that neural networks face challenges in directly classifying these components, which are nearly imperceptible to the human eye. However, NNs can be sensitive to the perturbation of high-frequency noise [34]. This is also an interesting topic of revealing the robustness of NNs, but it is beyond the scope of this article.

### v2

@@ Abstract
Gating mechanisms have emerged as an effective strategy integrated into model designs beyond recurrent neural networks for addressing long-range dependency problems. In a broad understanding, it provides adaptive control over the information flow while maintaining computational efficiency. However, there is a lack of theoretical analysis on how the gating mechanism works in neural networks. In this paper, inspired by the convolution theorem, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We investigate the interact between the element-wise product and activation functions in managing the responses to different frequency components. Leveraging these insights, we propose a Gating Mechanism Network (GmNet), a lightweight model designed to efficiently utilize the information of various frequency components. It minimizes the low-frequency bias present in existing lightweight models. GmNet achieves impressive performance in terms of both effectiveness and efficiency in the image classification task.
@@ 1 Introduction
The gating mechanism has been widely used in recurrent neural networks (RNNs) hochreiter1997long ; chung2014empirical to address long-term dependency issues. With the flourish of Transformers vaswani2017attention , self-attention-based architectures have largely replaced RNNs as the core model in natural language processing (NLP). Despite their success, scaling Transformers to long sequences remains challenging due to the quadratic complexity of global attention. Linear gating mechanisms (i.e. Gated Linear Units dauphin2017language ) have been rediscovered as an effective approach to address this problem.
Gated Linear Units (GLUs) introduce a gating mechanism that adaptively controls information flow, allowing the model to focus on relevant features and ignore less important ones. These architectures are simple to implement, and have no apparent computational drawbacks. Therefore, models built with GLUs have been developed in various tasks across NLP de2024griffin ; dubey2024llama ; gu2023mamba and Computer Vision liu2021pay ; ma2024rewrite ; yang2022focal . While architectures based on linear gating mechanisms have demonstrated the effectiveness, their theoretical analysis have not received enough attention. Except intuitive explanations, a deeper examination of their mathematical properties could provide valuable insights for the development of efficient gating mechanisms. Inspired by the Fourier interpretation of element-wise product, we propose to investigate training dynamics of neural networks (NNs) with gating mechanisms from a frequency perspective. This approach helps us understand how frequency components are selectively amplified or suppressed during learning. By the frequency analysis, we aim to uncover how GLUs influence the spectral properties of neural networks and affect the overall performance.
@@
In this paper, we systematically explore how gating mechanisms impact the training dynamics. We also provide a comprehensive understanding of how the element-wise product and activation functions interplay to shape the frequency response of neural networks. The first observation is inspired by the convolution theorem and convolution operation in the frequency domain. It indicates that element-wise multiplication in the spatial domain, equivalent to convolution in the frequency domain, inherently introduces new frequency components. It may enhance the learned representations for various frequency domains. However, as demonstrated in prior studies wang2020high ; yin2019fourier , NNs are sensitive to high-frequency noise, which may adversely influence training stability and generalization. Therefore, it is necessary to include activation functions within the gating units to regulate the newly introduced frequency components. Activation functions can broadly be classified based on their differential properties. Smooth activation functions (GELU hendrycks2016gaussian and Swish), characterized by continuous higher-order derivatives, tend to introduce fewer high-frequency components. On the other hand, activation functions (ReLU and ReLU6) with discontinuous higher-order derivatives are more likely to encourage the learning of high-frequency components.
@@
Benefited by the sufficient study from the frequency view, we find that the existing lightweight methods are not aware of how to efficiently and effectively utilize the high-frequency information. Capturing high-frequency components is especially challenging for efficient networks due to their limited capacity and reduced depth. These components often correspond to fine-grained details or subtle semantic distinctions critical for tasks such as fine-grained classification, edge detection, or video understanding. Without adequate modeling of high-frequency information, lightweight models may overly rely on coarse, low-frequency patterns, limiting their generalization and robustness. In this paper, we incorporate the simplest gated linear unit into the current block design and come up with a competitive lightweight model named as Gating Mechanism Network (GmNet). Our proposed network can more effectively capture information across different frequencies, offering both theoretical and practical advantages. With a simplified architecture of the network design, GmNet is able to achieve superior results compared with existing efficient modals without training with advanced techniques. For example, compared to EfficientFormer-L1 li2022efficientformer , GmNet-S3 outperforms $4.0\%$ on top-1 accuracy on ImageNet-1K with 4x faster latency on an A100 GPU.
@@
We succinctly summarize and emphasize the key contributions of this work as following points: 1). We present a comprehensive study on how gated linear units influence neural network training dynamics from a frequency perspective. Through both theoretical analysis and experimental evidences, we demonstrate how the interaction between element-wise multiplication and activation functions shapes the frequency response of NNs. 2).Inspired by the study of the gating mechanism, we investigate the effect of capturing the high-frequency information and propose a competitive lightweight model, named Gating Mechanism Network (GmNet), which achieves promising performance without the need for strong training strategy, surpassing numerous efficient designs.
@@ 2 Related Work
Gated Linear Units. Gating mechanisms have been extensively utilized in neural networks for sequence data processing tasks hochreiter1997long ; gers2000learning ; graves2005framewise ; dauphin2017language . With the rise of Transformer architectures, the Gated Linear Unit shazeer2020glu has been revisited and recognized as an efficient and effective enhancement for modern deep learning models. Recent research has increasingly focused on the integration of gating mechanisms into diverse network architectures of different tasks, including gate-MLP liu2021pay , Mamba gu2023mamba and Llama3 dubey2024llama . While many existing studies provide high-level insights into the functional role of gating mechanisms, there is a notable lack of in-depth analysis of their underlying learning processes. In this work, we conduct a series of experiments combined with solid theoretical analysis to elucidate the roles of gating mechanisms in enhancing the learning of neural networks. Specifically, we investigate the types of information that gating encourages networks to capture from a frequency perspective.
@@
Frequency Learning. Understanding the learning dynamics of neural networks from a frequency perspective is an increasingly interesting topic. It has been found that neural networks respond differently to frequency information depending on the tasks. It has been investigated that neural networks tend to learn low-frequency components quickly before fitting the high-frequency ones for regression tasks yin2019fourier ; tancik2020fourier ; rahaman2019spectral . In classification tasks, by investigating how neural networks respond to different frequency components during training, researchers found that there exists a trade-off between the accuracy and robustness wang2020high . NNs can improve the accuracy by learning more from high frequency components but they are also sensitive to high frequency noise which may disturb the training of NNs. In this paper, we investigate how GLUs optimize the trade-off which drive us to propose a new efficient model design.
@@
Lightweight Networks. The existing lightweight networks can be categorized into two types: (1) Pure convolution-based networks like MobileOne vasu2023mobileone , StarNetma2024rewrite and RepVitwang2024repvit ; (2) Self-attention-based architectures including EfficientModma2024efficient and EfficientFormerV2li2022efficientformer . Although existing network designs already achieve promising performance, they still exhibit significant shortcomings from the perspective of the frequency view. Both convolution and self-attention are learning methods with a low-frequency bias tang2022defects ; bai2022improving . Furthermore, because compact networks typically focus on minimizing parameters and simplifying model structure, their capacity to capture high-frequency details may be restricted. In this paper, we propose a new competitive and efficient network that demonstrates the effectiveness of learning from higher frequency components.
@@ 3 Revisiting Gating Mechanisms from A Frequency View
Figure 2: Block design of different variants of ResNet18 where $\odot$ represents the element-wise product and $\sigma$ means the activation function. In this paper, we investigate GLUs using ReLU6 and GELU as activation functions where corresponding variants can be named as Res18-Gate-ReLU6/GELU.
We begin by defining the components associated with different frequency bands and outlining the details of our experimental setup. With decomposing the raw data $\mathbf{z}$ into the high-frequency part $\mathbf{z}_{h}$ and the low-frequency part $\mathbf{z}_{l}$ where $\mathbf{z}=\mathbf{z}_{h}+\mathbf{z}_{l}$. Denoting a threshold $r$ and an image $\mathbf{x}$, we have the following equations:
$\mathbf{z}=\mathcal{F}(\mathbf{x}),\quad\mathbf{z}_{h},\mathbf{z}_{l}=\theta(\mathbf{z};R)$ (1)
where $\mathbf{z}=\mathcal{F}\left(\mathbf{x}\right)$ is the 2D Discrete Fourier Transform of $\mathbf{x}$. We select three vision backbones including ResNet-18 he2016deep , MobileNetv2 sandler2018mobilenetv2 and EfficientFormer-v2 li2023rethinking as representations to demonstrate the drawbacks of the CNN networks and transformer-based architectures on capturing the high frequency information and how the gating mechanism improve the capability of learning high-frequency components. Modifications to the network blocks of ResNet-18 are depicted in Fig. 2. We evaluate the classification performance on different frequency components of the input images at each training epoch. Changes in accuracy over time provide insights into the learning dynamics within the frequency domain wang2020high . To avoid the occasionality, we calculated the average over three training runs.
Figure 3: Comparison among Res18, Res18-Ewp. Res18-Gate-ReLU6 and Res18-Gate-GELU. We plot the learning curves of Resnet18 and its variants for $100$ epochs, together plotted with the accuracy of different frequency components $\mathbf{z}_{i}$. We set $r$ to $\{4,7,10\}$ respectively. All curves of $\mathbf{z}$ are from the test set. The legends can be found in the top of the figure. More results can be found in the supplementary material.
@@ 3.2 How Activation Function Works?
We begin with a brief mathematical analysis illustrating how the differential properties of activation functions influence the frequency characteristics of GLUs.
@@
where $f^{(n)}(t)$ represents the $n$-th derivative of $f(t)$. The proof of Lemma 3 can be easily obtained by the mathematical induction based on Theorem 2. From Eq. 3, we can find that $|F(\omega)|\leq\frac{C}{|\omega|^{n}}$ where $C$ is a sufficiently large constant. It indicates that if higher-order derivatives exist and are continuous, the high-frequency components of the Fourier transform decay more quickly with a rate of $|\omega|^{n}$. Conversely, if higher-order derivatives do not exist or are discontinuous, the decay of the high-frequency components of the Fourier transform slows down. Notably, for the smooth activation like $\operatorname{GELU}(x)=x\cdot\frac{1}{2}[1+\operatorname{erf}(x/\sqrt{2})]$, it is infinitely differentiable across the entire real domain, with all orders of derivatives being continuous and finite. The continuity and finite values of higher-order derivatives cause the high-frequency components of the Fourier transform to decay quickly, resulting in fewer high-frequency components.
To validate the above theoretical insights, we conduct experiments on ResNet18 and its variants to examine the impact of gating mechanisms and activation smoothness on frequency learning. As shown in Fig. 3, the testing curves of the model employing non-smooth activation functions (ReLU6) shows visibly surpass above the other models with different thresholds. It demonstrate superior performance in learning from high-frequency components. This supports the hypothesis that reduced smoothness in the activation function enhances the model’s ability to retain high-frequency information. Conversely, the variant using GELU exhibits stronger performance on low-frequency components, reflecting the strength of smooth activations in modeling low-frequency signals.
Figure 4: Comparison among different variants of MobileNetv2 and EfficientFormer-v2 (Eformerv2). Different architectures respond differently to specific frequency component. To ensure an informative comparison, we select representative frequency thresholds tailored to each model where we set $r$ to $\{6,10\}$ for MobileNetv2 and $\{4,8\}$ for EfficientFormer-v2 respectively. Additional results under other threshold configurations are included in the supplementary material.
4 Gating Mechanism Network (GmNet)
4.1 Rethinking Current Lightweight Model Architectures From A Frequency Perspective
@@
Figure 5: The illustration of GmNet architecture. GmNet adopts a traditional hybrid architecture, utilizing convolutional layers to down-sample the resolution and double the number of channels at each stage.
Before introducing our proposed network, we first investigate the importance of capturing high-frequency information in efficient architectures by modifying existing efficient models to incorporate Gated Linear Units (GLUs). Specifically, we select two representative architectures: the pure CNN-based MobileNetV2 sandler2018mobilenetv2 and the transformer-based EfficientFormer-V2 li2023rethinking . In both models, we replace the activation functions in their MLP blocks with a simple GLU. Detailed architectural modifications are provided in the supplementary material.
As shown in Fig. 4, we present the testing accuracy curves under varying frequency thresholds. Our results demonstrate that integrating GLUs improves classification accuracy on high-frequency components across both models, under different threshold settings. Notably, this improvement in high-frequency classification also correlates with a gain in overall performance. Furthermore, we observe that using GELU as the activation function within the GLUs enhances performance on low-frequency components, though it has a relatively minor effect on overall accuracy. These findings suggest that effectively modeling high-frequency information is more crucial for improving the performance of lightweight neural networks. It underscores the critical role of frequency-aware design in lightweight network architectures.
4.2 Architecture of GmNet
To address the limitation of low-frequency bias for current lightweight network designs, our proposed method named as GmNet integrates a simple gated linear unit into the block as illustrated in Fig. 5. GmNet offers both theoretical and practical advantages on encouraging the model to learn from a broader range of frequency regions, especially the high-frequency domain. We incorporate two depth-wise convolution layers with kernel sizes of $7\times 7$ at the beginning and end of the block respectively to facilitate the integration of low- and high-frequency information. At the core of the block, we have two $1\times 1$ convolution layers and a simple gated linear unit. We use the ReLU6 as the activation function. We vary the block numbers, input embedding channel numbers and channel expansion factors ‘ratio’ to build different sizes of GmNet, as detailed in Table 5.1.
5 Experiments
5.1 Results in Image Classification
Implementation details. We perform image classification experiments on the ImageNet-1K dataset, adopting a standard input resolution of $224\times 224$ for both training and evaluation. All model variants are trained from scratch for $300$ epochs using the AdamW optimizer, starting with an initial learning rate of $3\times 10^{-3}$ and a batch size of $2048$. The supplementary materials provide a comprehensive overview of the training setup. For performance assessment, we convert our PyTorch models into the ONNX format to measure latency on a Mobile device (iPhone 14) and a GPU (A100). Additionally, we deploy the models on the mobile device via CoreML-Tools to further evaluate latency. Importantly, our training approach does not incorporate advanced techniques such as re-parameterization or knowledge distillation. The results presented in Table 5.1 correspond to models trained without these enhancements.
Compared with the state-of-the-art. The experimental results are presented in Table 5.1. Without any strong training strategy, GmNet delivers impressive performance compared to many state-of-the-art lightweight models. With a comparable latency on GPU, GmNet-S1 outperforms MobileV2-1.0 by $3.5\%$. Notably, GmNet-S2 achieves $78.3\%$ with only $1.9$ms on the A100 which is a remarkable achievements for the models under 1G FLOPS. GmNet-S3 outperforms RepViT-M1.0 and StarNet-S4 by $1.9\%$ and $0.9\%$ in top-1 accuracy with $1.1$ ms and $1.4$ ms faster on the GPU latency, respectively. The improvements on the speed are over $30\%$. Additionally, with similar latency, GmNet-S3 delivers a $1.7\%$ improvement on the accuracy over MobileOne-S4. GmNet-S4 achieves 2x faster compared to RepViT-M1.5 on the GPU and it surpasses MobileOne-S4 of $2.1\%$ under the similar latencies of both GPU and Mobile. LeViT-256 graham2021levit matches the accuracy of GmNet-S4 but runs twice as slow on a GPU and 16 times slower on an iPhone 14 The strong performance of GmNet can be largely attributed to the clear insights of gating mechanisms and simplest architectures. Fig. 6 further illustrates the latency-accuracy trade-off across different models. More comparisons and results of downstream tasks including objective detection and segmentation can be found in supplementary.